# ReMA: Learning to meta-think for LLMs with multi-agent reinforcement learning

### v2

@@ Appendix H Supervised fine-tuning data collection
For experiments in Sec 4.4.1, we collect expert data to enhance the reasoning pattern, i.e. RL from SFT. Specifically, we collect demonstration data from GPT-4o Mini on MATH training dataset (7.5k problems) Hendrycks et al. (2021) and use it to fine-tune the LLMs. The data generation follows these steps: First, we prompt GPT-4o Mini to produce metacognitive reasoning for high-level model training. Specifically, we use different prompts to instruct it to rewrite and decompose a given question without providing a final answer. We collect metacognitive reasoning using two predefined actions, “rewrite” and “decompose”, which align with human approaches to complex problem-solving while preserving answer diversity. Next, we use the generated instructions to prompt GPT-4o Mini to follow the metacognitive steps and solve the question, obtaining SFT data for low-level policy training. Below, we present the prompts used for both high-level and low-level models.
@@
Appendix I Training Details
We refer to Appendix K for prompts we use during training. We implement the training pipeline with OpenRLHF (Hu et al., 2024a) which is a highly efficient codebase and is easy to scale up. We select REINFORCE++ to save resources and for efficient training. All experiments are conducted in a node of 8 NVIDIA A100 GPUs. We use bf16, Zero2, Flash-Attention and gradient checkpointing to run our experiments.
I.1 Training on MATH
VRP
MRP
MAMRP
I.2 Training on Reward Bench
VRP
MRP
MAMRP
Appendix J Qualitative Results
J.1 High-level policy finds better plans
J.2 Case study for Experiments in Sec 3.3
Fig 8 and Fig 9 show an case study of experiments in Sec 3.3.
Figure 8: Case Study for consistency reward of high-level agent
Figure 9: Case Study for base correctness reward of high-level agent
J.3 Case study for Experiments in Sec 4.5
Fig 10 shows the detailed outputs of the two groups of LLMs trained in JSON action. The outputs finally converge to relative easier and shorter sentences, compared to our main experiments. This may due to insufficient capacity of a small LM to produce a valid JSON-format response while simultaneously exploring diverse reasoning strategies.
Figure 10: Case Study for interpretability experiments in Sec 4.5
Appendix K Prompts
K.1 Prompts for Math problems
@@ I.2 Training on Reward Bench
MAMRP prompt: high-level agent:
@@
K.2 Prompts for LLM-as-a-Judge problems
We adopt the prompts from Saha et al. (2025a). VRP prompt:

### v3

@@ D.1.1 Supervised fine-tuning data collection
For experiments in Sec. 4.2.1, we collect expert data to enhance the reasoning pattern, i.e. RL from SFT. Specifically, we collect demonstration data from GPT-4o Mini on MATH training dataset (7.5k problems) Hendrycks et al. [2021] and use it to fine-tune the LLMs. The data generation follows these steps: First, we prompt GPT-4o Mini to produce metacognitive reasoning for high-level model training. Specifically, we use different prompts to instruct it to rewrite and decompose a given question without providing a final answer. We collect metacognitive reasoning using two predefined actions, “rewrite” and “decompose”, which align with human approaches to complex problem-solving while preserving answer diversity. Next, we use the generated instructions to prompt GPT-4o Mini to follow the metacognitive steps and solve the question, obtaining SFT data for low-level policy training. Below, we present the prompts used for both high-level and low-level models. Prompts can be found in Appendix G.1.1.
D.1.2 Dataset Curation of RewardBench970
[…]
JudgeBench 51.77 51.45 (-0.32) 50.65 (-1.12) 53.71 (+1.94)
Average 61.51 66.65 (+5.14) 65.53 (+4.02) 70.00 (+8.49)
JudgeBench 58.39 54.84 (-3.55) 55.81 (-2.58) 58.71 (+0.32)
Average 72.44 71.03 (-1.41) 68.06 (-4.38) 74.72 (+2.28)
D.1.3 Training on MATH
VRP
MRP
MAMRP
D.1.4 Training on Reward Bench
VRP
MRP
MAMRP
D.2 Multi-turn ReMA
We refer to Appendix G for prompts we use during training. We implement a multi-turn ReMA training pipeline with VeRL [Sheng et al., 2024] since it’s easier to implement complex training pipeline with a single centralized controller. Similar to OpenRLHF, VeRL is also a highly efficient and scalable codebase for further development.
For the multi-turn ReMA rollout, we use parameter sharing and simultaneous update by default. In details, we maintain two message lists with the system prompt of meta-thinking agent and reasoning agent respectively. During rollout, each agent act as ‘assistant’ in its own message list and the other agent act as ‘user’. We use three hyperparameters to control the rollout length: (1) ‘max_num_turns’: the maximum number of turns for each trajectory. (2) ‘max_response_length’: the maximum number of tokens for each turn’s response. (3) ‘max_prompt_length’: the maximum number of tokens for each trajectory.
During training, we apply the collected message list to Qwen2.5-7B’s chat template and build loss masks in order to compute the loss for all turns of one trajectory (message list).
Morever, for multi-turn ReMA rollout, unlike single agent single turn rollout, we need to carefully design the termination logic. Basically, we let the meta-thinking agent automatically decides when to finish the solving procedure, we use a special tag ‘[FINISH]’ to indicate the end of the solving procedure. After we detect this tag, we will terminate trajectory after the reasoning agent generates its output.
We also design other termination conditions to ensure the quality of the generated trajectories. If the last agent’s response is too long, we will terminate the whole trajectory and setting the reward to 0. We also introduce a different version of format reward: we give a reward of 1.0 only if the reasoning agent’s last turn response is correct and the meta-thinking agent’s last turn response include ‘[FINISH]’. We use math_verify as the default verifier.
D.2.1 SFT data collection of multi-turn MAMRP
We use GPT-4o to translate 817 samples in LIMO [Ye et al., 2025c] by prompting it to wrap each sentence with meta-thinking and reasoning tags. We use a temperature of 0. After filtering, we get 800 conversations for training. The prompt can be found in Appendix G.2.1. For supervised finetuning, we use LlamaFactory as the codebase and train the model for 3 epochs with a learning rate of 1e-5, consine learning rate scheduler, and batch size of 8. Use DeepSpeed Zero2 for distributed training.
D.2.2 Training on MATH
For training of multi-turn ReMA on MATH, we use GRPO [Shao et al., 2024] as the default learning algorithm. We refer to Appendix G.2.2 for prompts. For experiment in Sec 4.3, we use sample 128 prompts, each with 16 trajectories. During training, we drop the KL loss term to improve the numerical stability. We use a learning rate of 1e-6, bfloat16 precision, FSDP backend for distributed training. We split the rollout data into 4 mini-batches for update. For the sake of numerical stability, we do pre-clip before computing the exponential of log_prob for a upperbound of 3.0.
For the main result in Fig 5, we test different rollout configurations with a max_prompt_length of 4096, training for 500 steps. We use 32 NVIDIA A800 GPUs, the longest training cost about 40 hours due to large scale validation per 10 steps.
For the ablation results in Fig 6, we use a tiny subset of MATH Level 3-5, training for 300 steps. Specifically, we sample 19 questions for every single type (133 instances in total). We use 8 NVIDIA A800 GPUs, the training cost about 30 hours
We test different rollout configurations:
(1) max_num_turns=30, max_response_length=256, max_prompt_length=4096 (2) max_num_turns=30, max_response_length=1024, max_prompt_length=3072
And for the experiment of separate parameter in multi-turn ReMA, we iteratively train each agent with the same configuration as above, but with a switch interval of 10 steps, starting from the meta-thinking agent.
Appendix E Other Experiments
E.1 Reward functions shape cross-agent behaviors
We also investigate the impact of different reward function designs on ReMA’s behavior. In addition to the base reward setting described in Appendix C.2, we evaluate a consistency-based reward function using Qwen2.5-7B-Instruct. This reward function is designed to encourage the high-level agent to generate more detailed guidance. Indeed, we observe that the high-level agent trained in this manner produces more detailed solution steps compared to the one trained with the basic correctness format reward. However, we also find that this approach often leads to jailbreak behavior, where the high-level agent tends to include the final answer within its output, compromising the intended hierarchical reasoning process.
This suggests that (1) meta-thinking can not only emerge and be reinforced in the high-level agent but also in the low-level agent. During reinforcement learning (RL) training, the two agents develop a novel problem-solving pattern characterized by a role reversal. (2) Consistency-based rewards promote a more self-corrective approach at the lower level, potentially disrupting the intended separation of roles between planning and execution. For a detailed case study, refer to Appendix F.2.
E.2 Detailed Training Curves on Different Datasets of Multi-turn ReMA
We show the detailed training curves of the multi-turn ReMA on different datasets in Fig. 8.
Figure 8: Detailed Training Curves on Different Datasets of Multi-turn ReMA
Appendix F Qualitative Results
F.1 High-level policy finds better plans
F.2 Case study for Experiments in Section E.1
Fig. 10 and Fig. 11 show an case study of experiments in Sec. E.1.
Figure 10: Case Study for consistency reward of high-level agent
Figure 11: Case Study for base correctness reward of high-level agent
F.3 Case study for Experiments in Section 4.2.2
Fig. 12 shows the detailed outputs of the two groups of LLMs trained in JSON action. The outputs finally converge to relative easier and shorter sentences, compared to our main experiments. This may due to insufficient capacity of a small LM to produce a valid JSON-format response while simultaneously exploring diverse reasoning strategies.
Figure 12: Case Study for interpretability experiments in Section 4.2.2
Appendix G Prompts
G.1 Single-turn ReMA prompts
G.1.1 Prompts for JSON data collection
@@
G.1.2 Prompts for Math problems
@@
MAMRP prompt:
high-level agent:
@@
G.1.3 Prompts for LLM-as-a-Judge problems
We adopt the prompts from Saha et al. [2025a].
VRP prompt:
@@ G.1.3 Prompts for LLM-as-a-Judge problems
G.2 Multi-turn ReMA prompts
G.2.1 SFT data collection of multi-turn MAMRP
System prompt: You are classifying reasoning process data into two types of thinking. You will be given a question-answer pair from a reasoning dataset. Your task is to split all words into two parts. These words are crucial for analyzing reasoning patterns, so do not skip any details. - **Meta-Thinking Agent (MTA):** Responsible for high-level thought processes. This includes planning, evaluating steps, expressing uncertainty, making observations, or setting goals. Avoid detailed calculations. The content should be enclosed in ‘<meta_thinking>‘ and ‘</meta_thinking>‘. - **Reasoning Agent (RA):** Responsible for detailed problem-solving steps, such as calculations, logical deductions, or breaking down a problem into subproblems. The content should be enclosed in ‘<reasoning>‘ and ‘</reasoning>‘. **Rules to follow:** 1. **Do not assign large chunks of text to a single type of thinking.** The reasoning process consists of small, nonlinear thinking steps, so alternate appropriately between Meta-Thinking and Reasoning steps. 2. **Keep the words from the original solution unmodified whenever possible.** Words like "Wait," "Hmm," "But," etc., typically indicate Meta-Thinking and should be preserved. 3. **When finalizing the answer:** - The **Meta-Thinking Agent (MTA)** must explicitly confirm the answer before completion and output ‘[FINISH]‘. - The **Reasoning Agent (RA)** should then provide the final answer in the correct format. 4. **Do not skip any reasoning steps, even if they seem redundant, incorrect or irrelevant** 5. **Do not modify or remove any part of the original reasoning process**, even if it seems redundant or repetitive. The goal is to **preserve the exact flow of thought** as it naturally occurs. 6. **Retain all expressions such as "Wait," "Hmm," "But wait," etc., exactly as they appear. These indicate important cognitive processes and should not be skipped or altered.** Here are examples for you: [Examples] ... User prompt: [Begin of Question] {question} [End of Question] [Begin of Solution] {solution} [End of Solution]
G.2.2 Prompt for math problems
Meta-Thinking Agent (MTA):
System prompt: You are a meta-think agent that represents human high-level think process, when solving a question, you will have a discussion with human, each time you think about what to do next: e.g. - Exploring multiple angles and approaches - Breaking down the solution into clear steps - Continuously reflecting on intermediate results honestly and adapt your strategy as you progress - Backtracking when necessary - Requesting exploration of multiple solutions individually - Finally confirm the answer with the tag [FINISH] User prompt: {question}
Reasoning Agent (RA):
[…]